# Improving Autoformalization Using Direct Dependency Retrieval

### v1

@@ 4.3 Retrieval Results
Table 2 presents the dependency retrieval performance, evaluated on 200 samples across datasets of varying difficulty. The performance is measured by the average Precision, Recall, and F1-Score. Here, R@5 and R@10 denote the results of the RA method when retrieving the top-5 and top-10 most relevant dependencies, respectively.
The R@5 baseline exhibits suboptimal overall performance. As indicated in Table 1, the average dependency length across all difficulty levels is less than 5, which accounts for its low Precision. However, the concurrently low Recall suggests that the retrieved dependencies are inaccurate, leading to poor overall retrieval quality. By increasing the number of retrieved items to 10, R@10 achieves a higher Recall than R@5, but at the cost of a significant drop in Precision, resulting in a lower F1-Score. This phenomenon indicates that the traditional RAG method is not well-suited for autoformalization problems. Furthermore, Figure 3 illustrates the standard deviations of these metrics on 200 samples. The data reveals that the output quality stability, as measured by standard deviation, is comparable among DDR, R@5, and R@10. Nevertheless, DDR consistently and significantly outperforms existing RAG methods across all datasets in terms of the mean of Precision, Recall, and F1-Score. This result highlights the superiority of our proposed DDR approach.
@@ 4.4 Autoformalization Results
Table 3 reveals a complex relationship between model performance and dataset characteristics, with several findings that are counter-intuitive.
Figure 4: LLM consistency across multiple attempts. Subplots (a–e) correspond to datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot represents the number of samples ($y$ axis) whose outputs pass the BEq criterion at least $k$ ($x$ axis) times out of 8 attempts. Blue: RA+DDR; Orange: RA+R@10; Green: RAW
Figure 5: LLM efficiency across multiple attempts. Subplots (a–e) correspond to datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot shows the distribution of samples ($y$ axis) by the number of attempts ($x$ axis) required to pass the BEq criterion for the first time over 8 attempts. Blue: RA+DDR; Orange: RA+R@10; Green: RA
On the Diff01 subset, where inputs resemble unstructured natural language, SFT models generally exhibited poor performance. In contrast, ICL models, which are better adapted to general linguistic expressions, performed more effectively. Notably, on the Diff45 subset, which features a low average Depend Rate and Depend Length as shown in Table 1, SFT models demonstrated a performance improvement despite the higher difficulty level. ICL models also showed an anomalous increase in performance on this subset, contradicting the general trend where performance degrades as difficulty increases. For the other subsets, the results consistently show that as the difficulty level ascends, both the average Depend Rate and Depend Length within the dataset increase, corresponding directly with a steady decline in performance across all tested models.
Collectively, these observations establish a strong negative correlation between the performance of autoformalization models and the complexity of library dependency, as quantified by Depend Rate and Depend Length. This finding underscores the necessity of employing techniques such as RAG to aid models in managing such dependencies.
Against this backdrop, our proposed RA+DDR method demonstrates superior performance. As shown in Table 3, our method outperforms all baseline models on the BEq metric across nearly all subsets, including state-of-the-art models with tens of times more parameters. This result strongly indicates that autoformalization is a specialized challenge that requires targeted solutions beyond simple prompt engineering, further validating the value of our in-depth research into RAG.
4.5 Ablation Studies
Figure 4 evaluates the ablation methods from the perspective of multi-attempt reliability. It plots the number of problems for which the model’s generated Lean 4 statements were equivalent to the ground truth for n times (where n ranges from 1 to 8) across 8 independent generation attempts. The results indicate that, compared to the baseline methods RA and RA+R@5, our proposed method, RA+DDR, achieves higher consistency on a larger set of problems. RA+DDR outperforms the baselines, demonstrating the superior stability and reliability of its statement formalization results.
Figure 5 presents a comparison from the perspective of generation efficiency. It illustrates the distribution of problems based on the number of attempts required to achieve the first successful match with the ground truth across 8 independent trials. The experimental results show that the RA+DDR method tends to find the correct statement formalization in fewer attempts. In contrast, the success distribution for the baseline methods, RA and RA+R@5, is skewed towards requiring more trials to achieve the first match. This finding further highlights the efficiency advantage of our proposed method for the autoformalization task.
5 Discussion
DDR departs from traditional retrieval-based methods for formal dependency acquisition by employing an innovative generation-plus-verification paradigm. A key limitation of DDR, however, is its inability to precisely locate the generated items within the library. Specifically, whereas the full retrieval results in methods like liu2025rethinking typically include complete metadata such as the item’s name, code usage, and documentation, DDR’s potential to generate abbreviated or shortcut names precludes the retrieval of this supplementary information. This discrepancy renders DDR not directly compatible with the retrieval-augmented autoformalizer framework in liu2025rethinking. To accommodate this, during our inference experiments, we set all required information fields to None, apart from the dependency item names generated by DDR. Despite this incompatibility, experimental results demonstrate that DDR still outperforms traditional retrieval methods. The primary contribution of DDR lies in offering an innovative and flexible solution for the field of autoformalization. It introduces a more direct and efficient approach to dependency retrieval and enables the flexible integration of various general-purpose LLMs as the core of RAG frameworks.
@@ 6 Conclusion
We identify three core limitations of dependency retrieval methods in current autoformalization research: low retrieval accuracy; a reliance on a pre-defined hyperparameter for the number of dependencies; and an inability to learn and improve from ever-growing datasets. To address these challenges, we propose a novel RAG paradigm, termed DDR (Direct Dependency Retrieval). DDR leverages the contextual understanding capabilities of LLMs to directly generate candidate dependencies for a given informal statement. This is coupled with an efficient verification mechanism based on suffix arrays to ensure the validity of the generated outputs. This approach obviates the need for a pre-set number of dependencies and is designed to continuously improve its performance by training on expanding public datasets. Our experimental results demonstrate that DDR significantly outperforms traditional retrieval methods in terms of both recall and precision on the dependency retrieval task, and it exhibits a corresponding performance advantage in the downstream autoformalization task.
Appendix A Examples Illustration
@@
Figure 6: Examples from the FineLeanCorpus with difficulty levels 0 and 1. These problems are phrased in a more natural-language-like style rather than formal mathematical statements, which makes them more suitable for ICL-based models.
Appendix B Experiment Details

### v2

@@ 4.4 Retrieval Results
Although R@5 and R@10 are hallucination-free, their retrieval performance is questionable. To isolate retrieval quality from the issue of hallucination and enable a fair comparison of the methods’ true capabilities, we first filtered out all hallucinatory items from each method’s output and then recalculated their precision and recall. The performance comparison after filtering is shown in Table 3.
After filtering, the results show substantial performance disparities among the models. Despite having no hallucinations, R@5 and R@10 yield the lowest precision (0.02-0.10) and recall (0.10-0.20), indicating their retrieval capability is severely limited. The filtered prompt-based methods achieve moderate performance, with precision scores ranging from 0.2 to 0.5 and recall scores from 0.3 to 0.5. Our proposed DDR method demonstrates the most superior performance, achieving a high precision of 0.8-0.9 and a recall of 0.4-0.8, significantly outperforming all other baseline methods.
@@ 4.5 Autoformalization Results
The efficacy of DDR is further substantiated by its performance improvements in the downstream autoformalization task. As shown in Table 4 and Table 5, integrating DDR yields significant performance gains under both SFT and ICL frameworks. We specifically evaluated the combination of DDR with the RA model, an advanced model fine-tuned for autoformalization in liu2025rethinking. A format mismatch exists, as the RA model requires a comprehensive input format (full dependency name, informal description, and code), whereas DDR outputs only abbreviated dependency names. Consequently, the R@5 and R@10 retrievers from liu2025rethinking, designed for RA, are more format-compatible. Despite this mismatch, the combination of RA and DDR consistently outperforms all baseline methods. This success is primarily attributed to DDR’s superior retrieval precision and recall, highlighting its excellent generalization and robustness. In contrast, the performance of R@5 and R@10 reflects a trade-off between precision and recall, consistent with their respective retrieval characteristics.
To evaluate DDR’s potential as a plug-and-play component, we integrated it with ICL methods. Table 5 compares three experimental setups: (1) direct autoformalization, (2) two-stage autoformalization using a prompt-based dependency retrieval, and (3) two-stage autoformalization using a DDR-based dependency retrieval. The results indicate that the DDR-based RAG is the superior strategy in most test scenarios. Experiments on Deepseek-R1, illustrated in Figure 3, further reveal DDR’s stability advantage: the DDR-integrated method not only achieves a higher total pass@8 success count but also exhibits a significantly better distribution of successful attempts across 8 independent trials compared to the baseline. Furthermore, Figure 4 shows that this method can complete more tasks with fewer attempts.
@@ 5 Conclusion
We propose DDR, a method that leverages the contextual understanding of LLMs to directly generate candidate dependencies for a given informal statement in autoformalization tasks. This is complemented by SAC, which utilize suffix array for efficient dependency verification. Experimental results validate the effectiveness of our approach and demonstrate its potential in ICL hallucinations analysis and in assessing the difficulty of autoformalization datasets.
6 Limitations
A key limitation of DDR is its inability to precisely locate the generated items within the library. Specifically, whereas the full retrieval results in methods like liu2025rethinking typically include complete metadata such as the item’s name, code usage, and documentation, DDR’s potential to generate abbreviated or shortcut names precludes the retrieval of this supplementary information. This discrepancy renders DDR not directly compatible with the retrieval-augmented autoformalizer framework in liu2025rethinking.
DDR also exhibits limitations in handling certain special cases, which necessitate dedicated mechanisms. For instance, some core constructs (such as MOD), despite being crucial and frequently used in formalization, may not exist as standard library items
Formal mathematics libraries are continuously evolving, with dependency items being added, removed, or renamed across different versions. Such changes require corresponding maintenance of our system to ensure its continued functionality and effectiveness.
Appendix A Additional Results
Figure 5: LLM consistency across multiple attempts. From left to right correspond to Diff01, Diff23, Diff45, Diff67, and Diff89; from top to bottom correspond to Claude-3.5-Sonnet, DeepSeek-R1, GPT-4o, Ling-flash-2.0, and Qwen-Max. Each subplot illustrates the proportion of problems ($y$ axis) for which at least k (from 1 to 8, $x$ axis) out of 8 attempts successfully pass the BEq verification. Blue: DDR; Orange: ICL; Green: N/A.
Figure 6: LLM efficiency across multiple attempts. From left to right correspond to Diff01, Diff23, Diff45, Diff67, and Diff89; from top to bottom correspond to Claude-3.5-Sonnet, DeepSeek-R1, GPT-4o, Ling-flash-2.0, and Qwen-Max. Each subplot illustrates the intermediate pass@k score ($y$ axis) out of 8 attempts using BEq verification, k range from 1 to 8 ($x$ axis). Blue: DDR; Orange: ICL; Green: N/A.
While Figure 3 and Figure 4 primarily presented the experimental results on the Deepseek-R1 model, this section provides a more comprehensive evaluation by integrating the DDR method with various other ICL methods in Figure 5 and Figure 6. The experimental data demonstrates that incorporating DDR brings significant performance improvements to these baseline methods in the vast majority of cases.
Appendix B Examples Illustration of Diff01
A notable observation is that samples categorized as difficulty 0 and difficulty 1 present a counter-intuitively high formalization challenge. This is primarily because the natural language statements in this category often do not represent abstract descriptions of the problem’s mathematical essence, as illustrated by examples in Figure 7. Our subsequent experimental results corroborate this finding: for Diff01 dataset, the performance of autoformalization across all methods are worse than Diff23 and Diff45, yet still better than that of Diff89.
Figure 7 presents several examples from the Diff01 dataset. These problems are primarily formulated in a natural language style rather than as rigorous formal mathematical statements. Such presentation brings additional challenge in autoformalization, and renders difficulty metrics that rely on Depend Rate and Depend Length inapplicable, thereby impeding an accurate quantification of the autoformalization task difficulty on this dataset. Furthermore, because these problem descriptions lack an abstract representation of their mathematical essence, all methods evaluated in Table 4 and Table 5 exhibit a significant degradation in performance on the Diff01 dataset.
@@ Appendix B Examples Illustration of Diff01
Figure 7: Examples from the FineLeanCorpus with difficulty levels 0 and 1.
Appendix C Experiment Settings
We employ precision and recall as quantitative metrics to evaluate the performance of the dependency retrieval results. A key challenge in this evaluation arises from the discrepancy between the outputs of DDR, which directly generate dependency short names, and the ground-truth labels, which are Mathlib object identifiers extracted from source code. This leads to name ambiguity, as source code often contains unqualified or partially qualified names. To accurately identify true positives for our evaluation, we introduce a suffix-based matching criterion. Specifically, we represent each identifier as a sequence of dot-separated components. A retrieved identifier is considered a match for a ground-truth label if and only if two conditions are met: (1) the string representation of one identifier is a substring of the other; and (2) when both identifiers are split by the dot character (.) into lists of their components, one list is a sublist of the other.
To evaluate the semantic equivalence between the formal statements predicted by the model and the ground-truth formal statements, we employ two criteria: Type Checking (TypeC) and Bidirectional Extended Definitional Equality (BEq (liu2025rethinking). TypeC ensures the syntactic correctness of expressions, whereas BEq is utilized to determine their semantic equivalence. Building upon these criteria, we define the TypeC@8 and BEq@8 as our final evaluation metrics shown in Table 4 and Table 5. These metrics measure the percentage of problems for which at least one of the top-8 predicted statements is deemed equivalent to the ground-truth statement under the corresponding criterion.
To comprehensively evaluate our proposed method, we established several baseline models and conducted ablation studies. For the ablation studies, we employed the RAutoformalizer (RA) from liu2025rethinking as a pure formalizer baseline, and RA+R@k as an autoformalizer that incorporates the top-k retrieved dependencies using the method in liu2025rethinking. The RA was fine-tuned based on InternLM2-Math-Base-7B (ying2024internlm). Our baselines cover multiple aspects: first, we evaluated a range of state-of-the-art general LLMs via ICL, including DeepSeek-R1 (guo2025deepseek), Ling-flash-2.0 (ai2025ming), Claude-3.5-Sonnet (claude-3-5-sonnet-20241022) (anthropic_claude35_sonnet_2024), GPT-4o (gpt-4o-2024-08-06) (hurst2024gpt), and Qwen-Max (qwen-max-2024-09-19) (team2024qwen2); second, we included an InternLM2-Math-Base-7B (ying2024internlm) model fine-tuned on the MMA Lean subset (jiang2023multilingual). For our proposed DDR method, we selected Qwen3-32B (yang2025qwen3) as its base model. This choice highlights a core advantage of DDR: its flexibility in selecting a base dependency retrieval model without being restricted to specific embedding models.
During the decoding stage for all experiments, the temperature was set to 0.7, and we generated 8 samples for each problem. To ensure a fair comparison, the dependency retrieval step for the DDR method was performed only once, and its results were reused for all subsequent Lean 4 statement formalization generations. It is noteworthy that none of the autoformalization models discussed in the main text, including the fine-tuned ones, have undergone SFT on the FineLeanCorpus (peng2025criticlean).
@@ 5. Train Epochs: 6
Appendix D Prompt Templates
D.1 Prompt Template for Dependency Retrieval
You are an expert Lean 4 ‘Mathlib‘ dependency extractor.
Your task is to analyze a given mathematical proposition and identify the key mathematical concepts that would require specific definitions, theorems, or notations from the Lean 4 ‘Mathlib‘ library.
**CRITICAL INSTRUCTIONS:**
1. Your output **MUST** be a single, valid JSON object.
2. The JSON object **MUST** have one key: "dependency".
3. The value for the "dependency" key **MUST** be a list of strings. Each string is a potential ‘Mathlib‘ dependency.
4. The dependency names can be fully qualified (e.g., ‘Nat.minFac‘) or common shortcuts (e.g., ‘ncard‘, ‘Icc‘, ‘MOD‘).
5. If there is no potential ‘Mathlib‘ library dependency, output {"dependency":["No usage"]}
6. Do **NOT** write any Lean code. Do **NOT** provide any explanations or text outside of the JSON object. Your entire response must be only the JSON.
Here are some examples of the required input and output:
**Example 1**
**Math Proposition:**
[…]
1. f(n) is an integer for every integer n;
2. 2011^2 $\mid$ f(i) - a_i for i = 1, 2, $\ldots$, 2011;
3. 2011^2 $\mid$ f(n + 2011) - f(n) for every integer n;
is divided by 1000 is equal to 211.
**Output:**
dependency:
[Polynomial,degree,Fin,eval,Set.ncard]
**Example 2**
**Math Proposition:**
Prove that the set of integers (n $\ge$ 1) for which the smallest prime divisor of ((n!)^n + 1) is at most (n + 2015) is finite.
**Output:**
dependency:[Nat.factorial,Finite,
Nat.minFac]
**Example 3**
**Math Proposition:**
[…]
**Output:**
dependency:[MOD,Nat.Prime,Odd,Finset.Icc]
**Example 4**
**Math Proposition:**
Prove that if t represents the number ten, then the value of the expression (t + t + t + t) is 40.
**Output:**
dependency:[No usage]
Now, perform this task for the following proposition.
**Math Proposition:**
{input}
**Output:**
D.2 Prompt Template for Autoformalization with Dependency Retrieval
Please translate mathematical propositions into Lean 4 theorems. ‘Mathlib‘ is the MUST import.
DO NOT try to prove the theorem, ONLY translate it.
Use ‘sorry‘ as placeholder. Use the given potential ‘Mathlib‘ dependency items.
Here are some examples:
Math Proposition:
[…]
[Real,Continuous,Tendsto,Nat,atTop].
Lean Theorem:
import Mathlib
theorem exercise
(f : $\r Real$)
[…]
Math Proposition:
Let G be a finite group (with a multiplicative operation), and A be a subset of G that contains more than half of G’s elements. Prove that every element of G can be expressed as the product of two elements of A. The potential ‘Mathlib‘ dependency are
[Group,Finite,Set,ncard,Nat.card,Rat].
Lean Theorem:
import Mathlib
open Filter Topology
theorem exercise
[Group G]
(hG : Finite G)
(A : Set G)
(hA : A.ncard > (Nat.card G : Rat)/2):
[…]
Math Proposition:
[…]
[Nat,Nat.Prime,ZMod,ncard,MOD].
Lean Theorem:
import Mathlib
theorem exercise
(p : Nat)
(hp : Nat.Prime $\and p$ > 5)
[…]
Math Proposition:
Let p be 10, and q be 15, prove that p+b equals 25. The potential ‘Mathlib‘ dependency are
[No usage].
Lean Theorem:
import Mathlib
theorem exercise
(p : Nat)
(q : Nat)
(hp : p = 10)
(hq : q = 15):
p + q = 25 := sorry
Please translate the following proposition:
Math Proposition:
{input}
The potential ‘Mathlib‘ dependency are
{retrieved-dependency}
Lean Theorem: ‘‘‘
D.3 Other Templates
Other ICL templates employed in this experiment, including those for autoformalization without dependency retrieval and BEq verification, adhere to the corresponding settings in liu2025rethinking.